OpenStudy (anonymous):

If x = y + z and x + y =6 and x = 10, then z = _____?

OpenStudy (anonymous):

10 = y + z y has to be -4, because 10 + -4 = 6. 10 = -4 + z 10 = -4 + 14 So z = 14.

OpenStudy (anonymous):

You can click "Best Response" to give medals.